# Computational Complexity of Physical Information Sufficiency

### v3

@@ 2.8 Adjacent Objective Regimes and Bridge
  • horizon $T\in\mathbb{N}_{\geq 1}$ and history-dependent policy class,
@@
Definition 2.47 (Agent Snapshot vs Agent Process Typing).
@@
  • agent snapshot: fixed-parameter inference view (mapped to one-step deterministic class),
  • agent process: online/dynamical update views (horizon-extended, stochastic-criterion, or transition-coupled classes).
@@
In the represented adjacent family, direct transfer of static sufficiency claims is licensed if and only if the agent is typed as an agent snapshot. Every process-typed represented class has an explicit mechanized bridge-failure witness. (L: CC3, CC48, CC56.)
Operationally: a trained model evaluated with fixed parameters during inference is an agent snapshot for this typing; once parameters or transition-relevant internals are updated online, the object is process-typed and transfer from static sufficiency theorems is blocked unless the one-step bridge conditions are re-established.[C:P2.48;R:RA]
@@ 3.1 Structural Complexity vs Representational Hardness
In the mechanized architecture model, this is made explicit as a required-work observable and intrinsic lower bound: ‘requiredWork‘ is total realized work, ‘hardnessLowerBound‘ is the intrinsic floor, and ‘hardness_is_irreducible_required_work‘ proves the lower-bound relation for all nonzero use-site counts; ‘requiredWork_eq_affine_in_sites‘ and ‘workGrowthDegree‘ make the site-count growth law explicit.(L: DQ562,DQ560,DQ561,DQ563,DQ564,DQ565.)
@@ 3.2 Solver Integrity and Regime Competence
If $\mathsf{EXACT}$ and $\mathsf{EPSILON}(\varepsilon)$ have no evidence objects while $\mathsf{ABSTAIN}$ beats the inadmissibility branch, then RLFF strictly prefers $\mathsf{ABSTAIN}$ to both report types. (L: IC30, DQ570.)
@@
Hence $B_{\mathrm{raw}}(r)\leq B_{\mathrm{total}}(r)$ always, and strict inequality holds when evidence exists with positive certificate-bit allocation. (L: CC11, IC11, IC10, IC16, IC15, IC13, IC14, DQ567.)
@@
then there exists $n_{0}$ such that for all $n\geq n_{0}$, exact admissibility fails. The same conclusion holds in the canonical exponential-requirement specialization. (L: DQ573, DQ574.)
@@ 4 Computational Complexity of Decision-Relevant Uncertainty
This section establishes the computational complexity of information sufficiency, formalized as coordinate sufficiency, for an arbitrary decision problem $\mathcal{D}$ with factored state space. Because the problems below are parameterized by $\mathcal{D}$, and the $(A,S,U)$ tuple captures any problem with choices under structured information (Remark 1.1), the hardness results are universal: they apply to every problem with coordinate structure, not to a specific problem domain. We prove three main results:
@@ 4.6 Implications
4.7 The succinct–minimal gap and physical agents
@@
Let $C$ be a succinct circuit computing a utility function $U$, and let $C^{*}$ be the minimal circuit for the same function. The representation gap $\varepsilon=|C|-|C^{*}|$ measures how far the agent’s encoding is from the tightest possible encoding. This gap has no standard formal definition in the literature; it is the unnamed distance between an informal size class and a rigorous structural minimum.
@@
The results of this paper give $\varepsilon$ formal complexity consequences. By Theorem 4.7, determining the minimal sufficient coordinate set is coNP-complete under the succinct encoding. By Theorem 5.1, under ETH there exist succinct instances requiring $2^{\Omega(n)}$ time to verify sufficiency. In the formal Boolean-coordinate collapse model, size-bounded $\varepsilon=0$ search is exactly the relevance-cardinality predicate (Lean: S2P.min_representationGap_zero_iff_relevant_card). Together: for an agent encoded as a succinct circuit, the worst-case cost of determining whether its own coordinate attention set is minimal is exponential in the number of input coordinates (under ETH). The gap $\varepsilon$ is not merely unnamed; it is computationally irreducible: no polynomial-time algorithm solves all succinct instances (assuming $\mathrm{P}\neq\textsf{coNP}$), and the hard instance family requires $2^{\Omega(n)}$ time.
Physical agents as circuits.
Any physical agent that selects actions based on sensory state can be modeled, in the static snapshot, as a circuit: sensory inputs are coordinates $X_{1},\ldots,X_{n}$; the agent’s internal computation maps states to actions; and the quality of that mapping is measured by a utility function UU.[C:Prop. 2.48, Prop. 2.41;R:[bridge-admissible snapshot]] The explicit-state encoding (truth table) of this mapping has size exponential in the number of sensory dimensions. No physical agent can store it.[C:Thm. 5.1;R:[E] vs [S+ETH]] Every physical agent is therefore a succinct encoding, a circuit whose size is bounded by physical resources (neurons, synapses, transistors, parameters) rather than by the size of the state space it navigates.[C:Thm. 5.1;R:[S]]
This is not a metaphor. A biological nervous system is a feedforward or recurrent circuit that takes sensory state as input and produces motor commands as output.[C:Prop. 2.48;R:[represented adjacent class]] A trained neural network is a parameterized circuit.[C:Prop. 2.48;R:[represented adjacent class]] In both cases, the agent is a succinct encoding of a utility-to-action mapping, and the question “is this agent attending to the right inputs?” is exactly SUFFICIENCY-CHECK under the succinct encoding.[C:Thm. 7.2, Thm. 4.6;R:[S]] The compression that makes a circuit physically realizable, with polynomial size rather than exponential size, is the same compression that makes self-verification exponentially costly in the worst case: the state space the circuit compressed away is precisely the domain that must be exhaustively checked to certify sufficiency.[C:Thm. 5.1, Prop. 5.3;R:[S+ETH],[Q_fin]]
@@
The simplicity tax (Definition 9.2) measures $|\mathrm{Gap}(M,P)|=|R(P)\setminus A(M)|$: the number of decision-relevant coordinates that the agent’s model does not handle natively. In the $\varepsilon$ decomposition above, this is exactly the unattended-relevant component $|R_{\mathrm{rel}}\setminus I|$ (Lean: S2P.representationGap_missing_eq_gapCard, S2P.representationGap_eq_waste_plus_missing). For a physical agent modeled as a circuit, this is the coordinate-level manifestation of $\varepsilon$: the agent attends to a superset of what is necessary, or fails to attend to coordinates that are relevant, and faces worst-case cost exponential in $n$ to verify which case holds (Theorem 4.6). Theorem 9.3 then says the total relevance budget is conserved: unhandled coordinates are not eliminated, only redistributed as external per-site cost.
@@
Large language models and deep networks are succinct circuits with $\mathrm{poly}(n)$ parameters.[C:Thm. 5.1;R:[S]] Mechanistic interpretability asks which internal components (attention heads, neurons, layers) are necessary for a given capability; this is SUFFICIENCY-CHECK applied to the circuit’s internal coordinate structure. Pruning research asks for the minimal subcircuit that preserves performance; this is MINIMUM-SUFFICIENT-SET. The results of this paper imply that these questions are coNP-hard in the worst case under succinct encoding.[C:Thm. 4.6, Thm. 4.7;R:[S]] Empirical progress on interpretability and pruning therefore depends on exploiting the structural properties of specific trained circuits (analogous to the tractable subcases of Theorem 6.1), not on solving the general problem.[C:Thm. 4.6, Thm. 6.1;R:[S] vs [tractable regimes]]
@@
The formal results above apply to the static, deterministic model fixed by C1–C4 (Section 2.5). Biological neural systems add noise, temporal dynamics, and synaptic plasticity; trained networks add stochastic optimization and distribution shift. These factors may alter empirical tractability relative to the static worst-case bounds. The bridge conditions of Proposition 2.40 delineate when static sufficiency results transfer to sequential settings; beyond those conditions, the governing complexity objects change (Propositions 2.432.45). The claims in this subsection are therefore regime-typed consequences of the formal results, not universal assertions about all possible agent architectures.[C:Prop. 2.40, Thm. 2.46, Prop. 2.48;R:[represented adjacent class]]

### v4

@@ 2.8 Adjacent Objective Regimes and Bridge
  • horizon $T\in\mathbb{N}_{\geq 1}$ and history-dependent control-law class,
@@
Definition 2.55 (System Snapshot vs System Process Typing).
@@
  • system snapshot: fixed-parameter inference view (mapped to one-step deterministic class),
  • system process: online/dynamical update views (horizon-extended, stochastic-criterion, or transition-coupled classes).
@@
In the represented adjacent family, direct transfer of static sufficiency claims is licensed if and only if the system is typed as a system snapshot. Every process-typed represented class has an explicit mechanized bridge-failure witness. (L: CC3, CC48, CC56.)
Operationally: a fixed-parameter physical controller at runtime is a system snapshot for this typing; once parameters or transition-relevant internals are updated online, the object is process-typed and transfer from static sufficiency theorems is blocked unless the one-step bridge conditions are re-established.[C:P2.56;R:RA]
@@ 3.1 Structural Complexity vs Representational Hardness
Formally: this subsection separates class placement from regime-indexed access cost and states typed completeness for the declared static family.
@@
In the mechanized architecture model, this is made explicit as a required-work observable and intrinsic lower bound: ‘requiredWork‘ is total realized work, ‘hardnessLowerBound‘ is the intrinsic floor, and ‘hardness_is_irreducible_required_work‘ proves the lower-bound relation for all nonzero use-site counts; ‘requiredWork_eq_affine_in_sites‘ and ‘workGrowthDegree‘ make the site-count growth law explicit.(L: HD17, HD22, HD23, HD18, HD27, HD28.)
@@
Informally: the question stays fixed while representation changes cost, and the declared static regime set is fully covered.
@@ 3.2 Solver Integrity and Regime Competence
Formally: this subsection defines integrity, competence, typed report/evidence admissibility, and resource-bounded exact-claim limits.
@@
If $\mathsf{EXACT}$ and $\mathsf{EPSILON}(\varepsilon)$ have no evidence objects while $\mathsf{ABSTAIN}$ beats the inadmissibility branch, then RLFF strictly prefers $\mathsf{ABSTAIN}$ to both report types. (L: IC30, IC40.)
@@
Hence $B_{\mathrm{raw}}(r)\leq B_{\mathrm{total}}(r)$ always, and strict inequality holds when evidence exists with positive certificate-bit allocation. (L: CC11, IC11, IC10, IC16, IC15, IC13, IC14, IC18, IC19.)
@@
then there exists $n_{0}$ such that for all $n\geq n_{0}$, exact admissibility fails. The same conclusion holds in the canonical exponential-requirement specialization. (L: PBC7, PBC8.)
@@
Informally: honesty limits what you can claim, capability limits what you can solve, and hard regimes may require abstaining or using weaker exact claims.
@@ 4 Computational Complexity of Decision-Relevant Uncertainty
This section establishes the computational complexity of information sufficiency, formalized as coordinate sufficiency, for an arbitrary decision problem $\mathcal{D}$ with factored state space. Because the problems below are parameterized by $\mathcal{D}$, and the $(A,S,U)$ tuple captures any problem with choices under structured information (Definition 2.1), the hardness results are universal: they apply to every problem with coordinate structure, not to a specific problem domain. We prove three main results:
@@ 4.4 Anchor Sufficiency (Fixed Coordinates)
Remark 4.10 (Discovery vs Verification).
Informally, SUFFICIENCY-CHECK asks whether a given viewpoint works, a verification task ($\forall$), while ANCHOR-SUFFICIENCY asks whether any viewpoint works, a discovery task ($\exists\forall$). The proven complexity gap between coNP and $\Sigma_{2}^{\textsf{P}}$ suggests that finding the right frame is structurally harder than checking a given frame, unless the polynomial hierarchy collapses.
@@ 4.6 Implications
4.7 The succinct–minimal gap and physical controllers
@@
Let $C$ be a succinct circuit computing a utility function $U$, and let $C^{*}$ be the minimal circuit for the same function. The representation gap $\varepsilon=|C|-|C^{*}|$ measures how far the controller encoding is from the tightest possible encoding. This gap has no standard formal definition in the literature; it is the unnamed distance between an informal size class and a rigorous structural minimum.
@@
The results of this paper give $\varepsilon$ formal complexity consequences. By Theorem 4.7, determining the minimal sufficient coordinate set is coNP-complete under the succinct encoding. By Theorem 9.1, under ETH there exist succinct instances requiring $2^{\Omega(n)}$ time to verify sufficiency. In the formal Boolean-coordinate collapse model, size-bounded $\varepsilon=0$ search is exactly the relevance-cardinality predicate S2P2. Together: for a physical decision circuit encoded succinctly, the worst-case cost of determining whether its coordinate attention set is minimal is exponential in the number of input coordinates (under ETH). The gap $\varepsilon$ is computationally irreducible: no polynomial-time algorithm solves all succinct instances (assuming $\mathrm{P}\neq\textsf{coNP}$), and the hard instance family requires $2^{\Omega(n)}$ time.
Physical systems as circuits.
Any physical system that selects actions based on observed state can be modeled, in the static snapshot, as a circuit: observed inputs are coordinates $X_{1},\ldots,X_{n}$; internal computation maps states to actions; and the quality of that mapping is measured by a utility function UU.[C:Prop. 2.56, Prop. 2.49;R:[bridge-admissible snapshot]] The explicit-state encoding (truth table) of this mapping has size exponential in the number of sensory dimensions. No physical finite-memory system can store it.[C:Thm. 9.1;R:[E] vs [S+ETH]] Every realizable physical controller is therefore a succinct encoding, a circuit whose size is bounded by physical resources rather than by the size of the state space it navigates.[C:Thm. 9.1;R:[S]]
Atomic and molecular systems fit the same form: measured coordinates define observed state, transition operators implement action selection, and utility is objective-conditioned over resulting states.[C:Prop. 2.56;R:[represented adjacent class]] In this setting, the physical system is a succinct encoding of a utility-to-action mapping, and the question “is this system attending to the right coordinates?” is exactly SUFFICIENCY-CHECK under the succinct encoding.[C:Thm. 11.2, Thm. 4.6;R:[S]] The compression that makes a circuit physically realizable, with polynomial size rather than exponential size, is the same compression that makes self-verification exponentially costly in the worst case: the state space the circuit compressed away is precisely the domain that must be exhaustively checked to certify sufficiency.[C:Thm. 9.1, Prop. 9.3;R:[S+ETH], [Q_fin]]
@@
The simplicity tax (Definition 13.2) measures $|\mathrm{Gap}(M,P)|=|R(P)\setminus A(M)|$: the number of decision-relevant coordinates that the model does not handle natively. In the $\varepsilon$ decomposition above, this is exactly the unattended-relevant component $|R_{\mathrm{rel}}\setminus I|$ (S2P7, S2P5). For a physical circuit model, this is the coordinate-level manifestation of $\varepsilon$: the circuit attends to a superset of what is necessary, or fails to attend to relevant coordinates, and faces worst-case cost exponential in $n$ to verify which case holds (Theorem 4.6). Theorem 13.3 then says the total relevance budget is conserved: unhandled coordinates are not eliminated, only redistributed as external per-site cost.
@@
Large language models and deep networks are succinct circuits with $\mathrm{poly}(n)$ parameters.[C:Thm. 9.1;R:[S]] Mechanistic interpretability asks which internal components (attention heads, neurons, layers) are necessary for a given capability; this is SUFFICIENCY-CHECK applied to the circuit’s internal coordinate structure. Pruning research asks for the minimal subcircuit that preserves performance; this is MINIMUM-SUFFICIENT-SET. The results of this paper imply that these questions are coNP-hard in the worst case under succinct encoding.[C:Thm. 4.6, Thm. 4.7;R:[S]] Empirical progress on interpretability and pruning therefore depends on exploiting the structural properties of specific trained circuits (analogous to the tractable subcases of Theorem 10.1), not on solving the general problem.[C:Thm. 4.6, Thm. 10.1;R:[S] vs [tractable regimes]]
@@
The formal results above apply to the static, deterministic model fixed by C1–C4 (Section 2.5). Physical systems can add noise, temporal dynamics, and distribution shift. These factors may alter empirical tractability relative to the static worst-case bounds. The bridge conditions of Proposition 2.48 delineate when static sufficiency results transfer to sequential settings; beyond those conditions, the governing complexity objects change (Propositions 2.512.53). The claims in this subsection are therefore regime-typed consequences of the formal results, not universal assertions about all possible physical architectures.[C:Prop. 2.48, Thm. 2.54, Prop. 2.56;R:[represented adjacent class]]